### v1

# Entropic repulsion and scaling limit for a finite number of non-intersecting subcritical FK interfaces

Abstract
This article is devoted to the study of a finite system of long clusters of subcritical 2-dimensional FK-percolation with $q\geq 1$, conditioned on mutual avoidance. We show that the diffusive scaling limit of such a system is given by a system of Brownian bridges conditioned not to intersect: the so-called Brownian watermelon. Moreover, we give an estimate of the probability that two sets of $r$ points at distance $n$ of each other are connected by distinct clusters. As a byproduct, we obtain the asymptotics of the probability of the occurrence of a large finite cluster in a supercritical random-cluster model.
1 Introduction
Rigorous understanding of the behaviour of interfaces in statistical mechanics models has been the focus of intensive study for more than 50 years, especially in the case of the Ising model. The first rigorous results were perturbative and made use of the Pirogov–Sinaï theory to show that a low temperature two-dimensional Ising interface converges, after an appropriate diffusive scaling, towards a Brownian bridge [Gallavotti, higuchi]. However, these works are restricted to the very low temperature regime, even if the belief was that the result should hold for any subcritical temperature.
[…]
Later on, this technique has been found to be efficient for studying interfaces interacting with their environment. Indeed, the above mentioned works deal with unconstrained (also called free) interfaces, but recent works have been extending the study of these interfaces to broader settings in which non-trivial interactions with the environment are added. Let us cite [Ott_2018] for the case of a defect line in the Potts model, and — much more related to this work — [ottvelenikwachtelioffe] for the treatment of a Potts interface above a boundary wall. These examples of interfaces interacting with their close environment have turned out to be more delicate to handle and in certain conditions have been shown to exhibit highly non-trivial behaviours such as wetting transitions, which have been studied in [IV18].
Of the same nature is the study of a system of multiple interfaces interacting together. Such a problem will be the focus of this work. Indeed, this paper determines the scaling behaviour of a finite number of long clusters of subcritical Fortuin–Kasteleyn (FK) percolation, conditioned not to intersect. Regarding the discussion above, it can be seen as the study of a finite number of interfaces interacting together. An interesting feature of this setting is that when conditioned on not intersecting, the interaction between the clusters can turn out to be attractive, a priori allowing the existence of a pinning transition — a regime where this attraction is so strong that the clusters actually remain at a bounded distance from each other.
In this paper, we rule out the existence of such a transition. In the fashion of [ottvelenikwachtelioffe], we show that the behaviour of this system obeys an entropic repulsion phenomenon: the entropy caused by the large number of possible interfaces wins over the energetic reward obtained by staying close to each other, all the way down to the critical temperature. Such a phenomenon has been previously identified in a variety of settings, for instance in the three-dimensional semi-infinite Ising model at low temperatures [frohlichpfisterSemiInfiniteIsing], the 2+1-dimensional SOS model above a hard wall [caputomartinellitoninelliEntropicRepulsionSOS], a 1+1-dimensional interface above an attractive field in presence of a magnetic field [Ve2004] or a supercritical Potts interface above a wall [ottvelenikwachtelioffe], to mention but a few works studying this phenomenon.
In this work, entropic repulsion of the FK clusters at any subcritical temperature is established in Proposition LABEL:proposition_global_entropic_repulsion, which is probably the most important output of this work. As a byproduct, we derive two results regarding the global behaviour of such a system of conditioned interfaces. The first one is the diffusive scaling limit of such a system, which is shown to be a system of Brownian bridges conditioned not to intersect: the so-called Brownian watermelon. Moreover, we observe that the entropic repulsion phenomenon also allows the computation — up to a multiplicative constant — of the probability of the existence of such a system of interfaces. Finally, as a byproduct of the latter observation, we also obtain the asymptotics of the probability of the occurrence of a large finite connection in the supercritical random-cluster model.
The method is in spirit close to that of [ottvelenikwachtelioffe], but we face some additional difficulties — essentially due to the fact that the interaction is not only between a random interface and a deterministic object as in the latter work, but between several random interfaces, making the proofs more intricate as one has to control the joint behaviour of these interacting objects: in particular Section LABEL:section_RCM contains some estimates concerning the simultaneous behaviour of the system of clusters that appear to be new. The proofs crucially rely on the Ornstein–Zernike theory for subcritical random-cluster models, developed in [ozrandomclustercampanino].
1.1 Definitions of the random-cluster model and the Brownian watermelon
1.1.1 The random-cluster model
The model of interest is the so-called random-cluster model (also known as FK-percolation). We first recall its definition and a few basic properties (we refer to [duminilcopin2017lectures] for a complete exposition). The random-cluster model on $\mathbb{Z}^{2}$ is a model of random subgraphs of $\mathbb{Z}^{2}$. Its law is described by two parameters, $p\in[0,1]$ and $q>0$.
Let $G=\left(V(G),E(G)\right)$ be a finite subgraph of $\mathbb{Z}^{2}$. We denote its inner boundary (resp. outer boundary) by
$\partial G=\left\{x\in V(G),\exists y\notin V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}\\ (\text{resp. }\partial_{\mathsf{ext}}G=\left\{x\notin V(G),\exists y\in V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}).$
A percolation configuration on $G$ is an element $\omega\in\left\{0,1\right\}^{E(G)}$. We say that an edge $e\in G$ is open if $\omega(e)=1$ and closed otherwise. Two vertices $x,y\in\mathbb{Z}^{2}$ are said to be connected if there exists a path of nearest neighbour vertices $x=x_{0},x_{1},\dots,x_{n}=y$ such that the edges $\{x_{i},x_{i+1}\}$ are open for every $0\leq i\leq n-1$. In this case, we say that the event $\{x\leftrightarrow y\}$ occurs. A vertex cluster of $\omega$ is a maximal connected component of the set of vertices (it can be an isolated vertex). Given a percolation configuration $\omega$, we denote by $o(\omega)$ its number of open edges, and by $k(\omega)$ its number of vertex clusters.
A boundary condition on $G$ is a partition $\eta=P_{1}\cup\dots\cup P_{k}$ of $\partial G$. From a configuration $\omega\in\left\{0,1\right\}^{E(G)}$, we create a configuration $\omega^{\eta}$ by identifying the vertices that belong to the same $P_{i}$ of $\eta$. Two particular boundary conditions, that we shall call the free boundary condition (resp. wired boundary condition), consist in the partition made of singletons (resp. of the whole set $\partial G$). We shall write $\eta=0$ (resp. $\eta=1$) for this specific boundary condition.
Definition 1.1.
Let $G=\left(V(G),E(G)\right)$ be a finite subgraph of $\mathbb{Z}^{2}$, and $\eta$ be a boundary condition on $G$. Let $p\in[0,1]$ and $q>0$. The random-cluster measure on $G$ with boundary condition $\eta$ is the following probability measure on percolation configurations on $G$:
$\phi^{\eta}_{p,q,G}\left(\omega\right)=\frac{1}{Z^{\eta}_{p,q,G}}\left(\frac{p}{1-p}\right)^{o(\omega)}q^{k(\omega^{\eta})},$
where $Z^{\eta}_{p,q,G}>0$ is the normalisation constant ensuring that $\phi^{\eta}_{p,q,G}$ is indeed a probability measure. We shall refer to $Z^{\eta}_{p,q,G}$ as the partition function of the model.
It is classical that for $\eta=0$ and $\eta=1$, this measure can be extended to the whole plane $\mathbb{Z}^{2}$, by taking the weak limit of the measures $\phi^{\eta}_{p,q,G_{n}}$ over any exhaustion $\left(G_{n}\right)_{n\in\mathbb{N}}$ of $\mathbb{Z}^{2}$, and that the limit measure does not depend of the choice of the exhaustion. Below, we will simply write $\phi_{p,q}^{\eta}$ instead of $\phi_{p,q,\mathbb{Z}^{2}}^{\eta}$.
A very fundamental feature of this model is that it undergoes a phase transition. Namely for any $q\geq 1$, there exists a critical parameter $p_{c}=p_{c}(q)\in(0,1)$ such that:
  • $\forall p<p_{c}(q),\phi^{1}_{p,q}\left(0\leftrightarrow\infty\right)=0$;
  • $\forall p>p_{c}(q),\phi^{0}_{p,q}\left(0\leftrightarrow\infty\right)>0$,
where $\{0\leftrightarrow\infty\}$ is the event that the cluster of 0 is infinite.
We are going to be interested in the first case — called the subcritical regime. In this case it is well known that the choice of boundary conditions does not affect the infinite volume measure. We thus drop $\eta$ from the notation and simply write $\phi_{p,q}$ for the unique infinite volume measure when $p<p_{c}(q)$. Another important feature of the subcritical random-cluster model is the existence and the positivity of the following limit:
$\tau_{p,q}:=\lim_{n\rightarrow\infty}-\frac{1}{n}\log\left[\phi_{p,q}\left(0\leftrightarrow(n,0)\right)\right].{\hbox{\rm(1)}\kern 5.69046pt\kern-5.69046pt}eftau{=}D0{1}$

### v2

@@
 These examples of interfaces interacting with their close environment have turned out to be more delicate to handle and in certain conditions have been shown to exhibit highly non-trivial behaviours such as wetting transitions, which have been studied in [IV18].
-Of the same nature is the study of a system of multiple interfaces interacting together.
-Such a problem will be the focus of this work.
-Indeed, this paper determines the scaling behaviour of a finite number of long clusters of subcritical Fortuin–Kasteleyn (FK) percolation, conditioned not to intersect.
-Regarding the discussion above, it can be seen as the study of a finite number of interfaces interacting together.
-An interesting feature of this setting is that when conditioned on not intersecting, the interaction between the clusters can turn out to be attractive, a priori allowing the existence of a pinning transition — a regime where this attraction is so strong that the clusters actually remain at a bounded distance from each other.
-In this paper, we rule out the existence of such a transition.
-In the fashion of [ottvelenikwachtelioffe], we show that the behaviour of this system obeys an entropic repulsion phenomenon: the entropy caused by the large number of possible interfaces wins over the energetic reward obtained by staying close to each other, all the way down to the critical temperature.
+Of the same nature is the study of a system of multiple interacting interfaces, which is the focus of the present work.
+Indeed, this paper determines the scaling behaviour of a finite number of long clusters of subcritical Fortuin–Kasteleyn (FK) percolation, conditioned not to intersect; subcritical percolation clusters mimic interfaces in the low temperature regime.
+An interesting feature of this setting is that when conditioned not to intersect, the interaction between the clusters can turn out to be attractive, a priori allowing the existence of a pinning transition — a regime where this attraction is so strong that the clusters actually remain at a bounded distance from each other.
+We rule out the existence of such a transition.
+In the fashion of [ottvelenikwachtelioffe], we show that the behaviour of this system obeys an entropic repulsion phenomenon: the entropy caused by the large number of possible clusters wins over the energetic reward obtained by keeping them close together, all the way up to the critical point.
 Such a phenomenon has been previously identified in a variety of settings, for instance in the three-dimensional semi-infinite Ising model at low temperatures [frohlichpfisterSemiInfiniteIsing], the 2+1-dimensional SOS model above a hard wall [caputomartinellitoninelliEntropicRepulsionSOS], a 1+1-dimensional interface above an attractive field in presence of a magnetic field [Ve2004] or a supercritical Potts interface above a wall [ottvelenikwachtelioffe], to mention but a few works studying this phenomenon.
 In this work, entropic repulsion of the FK clusters at any subcritical temperature is established in Proposition LABEL:proposition_global_entropic_repulsion, which is probably the most important output of this work.
-As a byproduct, we derive two results regarding the global behaviour of such a system of conditioned interfaces.
+As a byproduct, we derive two results regarding the global behaviour of such a system of conditioned clusters.
 The first one is the diffusive scaling limit of such a system, which is shown to be a system of Brownian bridges conditioned not to intersect: the so-called Brownian watermelon.
-Moreover, we observe that the entropic repulsion phenomenon also allows the computation — up to a multiplicative constant — of the probability of the existence of such a system of interfaces.
+Moreover, we observe that the entropic repulsion phenomenon also allows the computation — up to a multiplicative constant — of the probability of the existence of such a system of clusters.
 Finally, as a byproduct of the latter observation, we also obtain the asymptotics of the probability of the occurrence of a large finite connection in the supercritical random-cluster model.
-The method is in spirit close to that of [ottvelenikwachtelioffe], but we face some additional difficulties — essentially due to the fact that the interaction is not only between a random interface and a deterministic object as in the latter work, but between several random interfaces, making the proofs more intricate as one has to control the joint behaviour of these interacting objects: in particular Section LABEL:section_RCM contains some estimates concerning the simultaneous behaviour of the system of clusters that appear to be new.
-The proofs crucially rely on the Ornstein–Zernike theory for subcritical random-cluster models, developed in [ozrandomclustercampanino].
+The method is in spirit close to that of [ottvelenikwachtelioffe], but with considerable additional difficulties.
+These are essentially due to the fact that the interaction is not only between a random object and a deterministic one, but between several random objects, forcing one to control their joint behaviour.
+The proofs make heavy use of the Ornstein–Zernike theory for subcritical random-cluster models, developed in [ozrandomclustercampanino].
 1.1 Definitions of the random-cluster model and the Brownian watermelon
@@
 Let $G=\left(V(G),E(G)\right)$ be a finite subgraph of $\mathbb{Z}^{2}$. We denote its inner boundary (resp. outer boundary) by
-$\partial G=\left\{x\in V(G),\exists y\notin V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}\\ (\text{resp.
-}\partial_{\mathsf{ext}}G=\left\{x\notin V(G),\exists y\in V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}).$
+$\partial G$ $=\left\{x\in V(G),\exists y\notin V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}\text{ and}$
+$\partial_{\mathsf{ext}}G$ $=\left\{x\notin V(G),\exists y\in V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\},\text{ respectively.}$
 A percolation configuration on $G$ is an element $\omega\in\left\{0,1\right\}^{E(G)}$.